OpenStudy (anonymous):

how would i solve (x double dot) + x(x dot)+x=0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

help

OpenStudy (anonymous):

I think you can solve it by reduction of order

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

you deleted that? it wasnt the same

OpenStudy (anonymous):

yea, i had it wrong

OpenStudy (anonymous):

that one can be solved differently

OpenStudy (anonymous):

i thought reduction of order was for coefficients which are random polynomials of in t

OpenStudy (anonymous):

\[x \prime \prime+xx \prime+x=0\]

OpenStudy (anonymous):

for this you say let u=dx/dt than d^2x/dt^2=u du/dt

OpenStudy (anonymous):

than the equation will be u du/dt + xu+x=0

OpenStudy (anonymous):

ahh, that is clever, i will try that

OpenStudy (anonymous):

thanks a ton

OpenStudy (anonymous):

and I guess you can use an integrating factor to solve this

OpenStudy (anonymous):

now I have to go to the shop. If you have a problem just post it

OpenStudy (anonymous):

sweet, thanks a ton you are amazing

OpenStudy (anonymous):

that integral doesn't make sense, it is with respect to the wrong variable

OpenStudy (anonymous):

you have to change u back to x

OpenStudy (anonymous):

I did not solve it, what did you get?

OpenStudy (anonymous):

ah, i'm saying even setting up the integration factor

OpenStudy (anonymous):

dot means that we differentiate with respect to t

OpenStudy (anonymous):

x is a function of t, but the integral should be with respect to t

OpenStudy (anonymous):

yea, i set up \[u \prime+ux+x=0\]

OpenStudy (anonymous):

u du/dt + xu+x=0 du/dt +u/x=-x

OpenStudy (anonymous):

u'=u du/dt by the chain rule

OpenStudy (anonymous):

e^integral1/x is the integrating factor that is just x

OpenStudy (anonymous):

so multiplying with x gives xdu/dt+u=-x^2

OpenStudy (anonymous):

(xu)dot=-x^2

OpenStudy (anonymous):

xu=-(x^3)/3 +C

OpenStudy (anonymous):

u=-(x^2)/3 +c/x

OpenStudy (anonymous):

u=dx/dt

OpenStudy (anonymous):

dx/dt=-(x^2)/3 +c/x

OpenStudy (anonymous):

so t= -(x^3)/9 +Clnx

OpenStudy (anonymous):

but when you set up that equation du/dt +u/x=-x we should be integrating with respect to t

OpenStudy (anonymous):

you are integrating with respect to x

OpenStudy (anonymous):

for the integrating factor no

OpenStudy (anonymous):

it is only for making it a full derivative

OpenStudy (anonymous):

the integrating factor makes sense

OpenStudy (anonymous):

I think it is like this, but I would not bet on it :-)

OpenStudy (anonymous):

the actual integratioin though does not and why is the equation u du/dt rather than du/dt

OpenStudy (anonymous):

u=dx/dt u dot= x double dot

OpenStudy (anonymous):

this part is not 100% clear for me :) but I am sure that is right and comes from the chain rule

OpenStudy (anonymous):

I just posted a question about it, maybe some1 is wise

OpenStudy (anonymous):

(I'm just a 1st year maths student)

OpenStudy (anonymous):

well, thank you for your help, i will check back later. i really appreciate it andras

OpenStudy (anonymous):

your welcome